OpenStudy (anonymous):

I need to find the integral from -6 to 0 for sqrt (36-x^2). Have been trying for 20 minutes and can't get right answer. Need to know how to start because obviously I am not starting it right.

OpenStudy (anonymous):

Take a substitution of x=6sin(t) I think u can carry rest..if not tell me

OpenStudy (anonymous):

Where are you getting the sin from?

OpenStudy (anonymous):

\[\int\limits_{-6}^{0}\sqrt{(36-x^{2}})dx\]

OpenStudy (anonymous):

\[\int\limits_{-6}^{0}\sqrt{36-x^2}dx\]Let x=6sint then, dx=6cost dt

OpenStudy (anonymous):

\[\int\limits_{-\pi/2}^{0}\sqrt{36-36\sin^2t}6\cos tdt\]

OpenStudy (anonymous):

now you have 36(1-sin^2 (t) in the root sign. sin^2(t)+cos^2(t)=1 to eliminate the root. go from there...

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+sqrt+%2836-x%5E2%29. Click on "Show steps"

OpenStudy (anonymous):

How do I mark that my question was given a good answer?

OpenStudy (anonymous):

click good answer on one of us to give a medal

OpenStudy (anonymous):

\[\int\limits_{-\Pi/2}^{0}\sqrt{6^2-6^2\sin^2(t)}\cos(t)dt\] \[\int\limits_{-\Pi/2}^{0}6\sqrt{1-\sin^2(t)}\cos(t)dt\] \[\int\limits_{-\Pi/2}^{0}6\sqrt{\cos^2t}\cos(t)dt\] \[\int\limits_{-\Pi/2}^{0}6{\cos(t)}\cos(t)dt\] \[\int\limits\limits_{-\Pi/2}^{0}{6\cos^2(t)dt}\] \[\int\limits\limits_{-\Pi/2}^{0}6{(\cos(2t)+1)/2}dt\] \[[\sin(2t)/2 +t](0,-\Pi/2)\] \[\Pi/2\]

OpenStudy (anonymous):

Where do I do that ? It says that I need to log in but it says that I am already logged in.

OpenStudy (anonymous):

weird. try leaving the site and coming back.

OpenStudy (anonymous):

Got it! Thanks for everyones help. I see what I did wrong now.

OpenStudy (anonymous):

Medals given!!